What is the slope of the line 2x + 2y = -7?
Answer by nyc_function(2741)

The slope is the coefficient of x in the equation y = mx + b.

I need to write your equation in the general form above.

To do so, I will need to ISOLATE y.

2x + 2y = -7

2y = -2x - 7

y = -2x/2 -7/2

y = -x - 7/2

Do you see it now?

Your slope is -1.